OpenStudy (anonymous):

HELP!GIVING OUT MEDALS! Rename 0.245 as a percent A.245% B.24.5% C.2.45% D.25% Which of the following is greater than 0.366? Put 0.367, 36.5%, and 9/25 in order from least to greatest. A. 9/25 , 0.367, 36.5% B.9/25, 36.5%, 0.367 C.0.367,9/25,36.5% D.36.5%,0.367,9/25 On a city block, 4/5 of the people subscribe to the newspaper. What percent subscribe to the newspaper? A.60% B.80% C.40% D.20%

undeadknight26 (undeadknight26):

First one is D because you would round it to the nearnest 10th place.

OpenStudy (anonymous):

3rd one is B

OpenStudy (anonymous):

2nd one is B also

OpenStudy (anonymous):

@undeadknight26 i got that wrong! its actually B >.< & thanks @CashBoy27

undeadknight26 (undeadknight26):

o.o Gomen...

OpenStudy (anonymous):

1st one is B. 0.245 in a percent is 24.5

OpenStudy (anonymous):

.367 36.5% = .365 9/25 = .36 from least to greatest = 9/25, 36.5%, .367 (answer B) ============ 4/5 = .80 = 80% (answer B)

OpenStudy (anonymous):

There all B